# NetGPT: An AI-Native Network Architecture for Provisioning Beyond Personalized Generative Services

### 18 Dec 2023 v3

@@ I Introduction
With the remarkable success of deep learning spanning from decision-making in AlphaGo to human-level interaction like ChatGPT, it is anticipated that artificial intelligence (AI) will be embodied in 6G networks. Along with the enhanced edge computing capabilities, AI could benefit effective orchestration of network resources and improve the quality of service (QoS). Correspondingly, investigation on efficient AI-based service provisioning has attracted an intense research interest. On the other hand, application of one AI model is often limited to certain scenarios or tasks. In this context, large language models (LLMs) (e.g., generative pre-trained transformer, GPT) could perform well in various natural language processing (NLP) and computer vision tasks. These inspiring advancements shed light on revolutionizing cellular networks by LLMs. For instance, [1] harnesses collective intelligence for efficient network management, by delving into the deployment of on-device LLMs and proposing a multi-agent system architecture. Similarly, [2] challenges traditional network paradigms for LLM training, and proposes a novel, cost-effective architecture tailored to LLM-specific communication patterns, with a demonstrated $75\%$ network cost reduction without sacrificing the performance. These progresses [1, 2] underscore the importance of integrating LLMs with innovative network architectures, a key to unlock greater efficiency and performance in advanced network environments. Notably, fine-tuning is still a prerequisite to align pre-trained LLMs to follow human intents [3] and yield personalized outputs. Therefore, it might be cost-ineffective to simply deploy multiple copies of bloated model parameters to support different purposes, and a feasible solution remains under-investigated.
@@ II Implementation Showcase of NetGPT
As illustrated in Fig. 2, we present a synergistic cloud-edge framework to accomplish personalized generative services, by leveraging distinctive pre-trained LLMs for cloud and edge (e.g., base stations [BSs]) deployment. In particular, limited by the availability of open-source LLMs, we select and deploy the LLaMA-7B model [8] and the GPT-2-base model, which consist of approximately $6.7$ and $0.1$ billion parameters, at the cloud and the edge, respectively. However, it should be noted that NetGPT allows the utilization of other LLMs as well. On this basis, we delve into implementation details of cloud-edge LLM synergy towards NetGPT in an incremental manner. In particular, we start with detailed DNN structures of two LLMs (i.e., LLaMA-7B model and GPT-2-base model). Then, we discuss the effective means to fine-tune these LLMs on computation-limited devices, and demonstrate the effectiveness for location-based personalized generative services after synergizing edge LLMs and cloud LLM. Notably, the “LLM synergy” framework significantly contrasts with split learning [5] and federated learning [3], which aims to train the divided segments of a DNN on different clients, or jointly learn from data distributed across multiple nodes without compromising the data privacy. Orthogonal to split learning and federated learning, our framework focuses on the effective cloud-edge collaboration with prompt enhancement & de-duplication at the edge and personalized responses at the cloud.
@@ II-B Low-Rank Adaptation and Cloud LLM Fine-Tuning
As LLaMA lacks the capability to generate responsive text [8], an extra fine-tuning is still required. However, a direct fine-tuning of LLMs such as a LLaMA still requires significant computational resources. For example, it demands $112$ GB video random access memory (VRAM) to fine-tune the LLaMA-7B model, far more than the capacity of NVIDIA A100 Tensor Core GPU. Therefore, we leverage a low-rank adaptation (LoRA) technique [11], a parameter efficient fine-tuning (PEFT) technique, to achieve parameter-efficient fine-tuning on a consumer-level hardware. Notably, for NetGPT, PEFT techniques are not indispensable components and LoRA only demonstrates the feasibility of fine-tuning on computation-limited network elements of NetGPT.
@@ II-C2 LLM-instructed data collection:
Different from the cloud LLM, which is fine-tuned with web-based conversational datasets [12], so to respond to “comprehensive prompts” in a manner that is informed, contextually rich, and aligns with the broader conversational context, the data collection for edge LLMs requires extra efforts. This differentiation in data sources underlines the collaborative efficacy of the synergy architecture, with each component playing a specialized role in the data processing workflow. In order to implement personalized edge LLM, it is crucial to grant the GPT-2-base model the capability to extend a “concise” prompt by appending location-based information. Basically, the positioning information can be conveniently obtained according to the locations of affiliated BSs stored in the 5G access and mobility management function (AMF). Meanwhile, in order to complement more comprehensive information, we take the self-instruct approach [13] and leverage OpenAI’s Text-Davinci-003 model to generate useful text samples. In particular, as for each location, we use a set of manually-written location-related prompts to interact with the OpenAI’s Text-Davinci-003 model, and leverage the generative response texts as the “intended prompt”, which augments the most likely word following the “concise prompt” and describes the corresponding more comprehensive intent. For example, as illustrated on the top-left side of Fig. 3, from a perspective of real-life linguistic patterns, “libraries” frequently correlates with a contextual word “collections”. Besides, the top-right part of Fig. 3 demonstrates how post-fine-tuning edge LLM can generate an elongated and more comprehensive prompt in response to the “concise prompt”. On this basis, a series of mappings between the “concise prompt” and an “intended prompt” can be collected. Considering the size and task complexity of the edge LLM, we collect a dataset comprising approximately $4,000$ samples for directly fine-tuning the GPT-2-base model towards a prompt-completion model. Notably, for scenarios where stronger generality is required, edge LLMs can be enhanced with a larger-scale LLM, and fine-tuning techniques such as LoRA can be employed as well.
@@ III AI-Native Network Architecture Towards NetGPT
We argue that NetGPT provides the opportunity to transform cellular networks into an AI-native networking architecture, which provisions personalized, networked and inclusive intelligence for end users and grants users more privilege to access generative services anytime and anywhere. Nevertheless, such a transformation does come at a cost. It requires substantial changes, far more than installing racks of servers at the edge location and local break-out of the traffic for edge processing. In particular, compared with conventional connectivity-oriented communications systems, wherein a typical service establishes connections between two specific terminals and the communication source and destination are clearly defined by end users, NetGPT requires to establish generative performance-driven connections in a more implicit manner. Moreover, as NetGPT involves more frequent data collection and processing modules for training personalized LLM models, computing resources shall be consistently scheduled to accomplish NetGPT. In other words, as shown in Fig. 4, NetGPT necessitates the design of deeply converged C&C in radio access networks (RANs). On top of these novel features, a logical AI workflow shall be devised to establish (beyond) generative service orchestration.
III-A Converged C&C in RAN
In order to effectively organize heterogeneous resources, which may simultaneously cover terrestrial and non-terrestrial communications, the RAN for NetGPT has to first provide a control plane (CP) for seamless connectivity control to ubiquitously support prompting and generative information transmission in the user plane (UP). Such elements can be developed in accordance with the 5G and 5G-Advanced techniques. In addition, it is worthwhile to introduce an independent computing plane (CmP) to coordinate computing resources and perform AI-related functionalities, so as to facilitate the deployment and update of generative services.
@@ III-C Personalized Profiling
III-D C&C Resource Management
As part of provisioned services in future cellular networks, the orchestration of resources for NetGPT shares some similarities as that for other network services, including connectivity establishment and computing resource allocation. However, it also poses additional challenges, since the scope of resources spans distributed nodes from the cloud to the terminal. Therefore, novel protocol stack needs be carried on radio control signaling (e.g., RRC) or radio data protocol (e.g. PDCP or SDAP) to transmit AI-generative messages and implement model update & distribution.
III-E Logical AI Workflow
@@ V Conclusion
  • Besides network optimization at higher layers, how to develop LLM-based physical and radio link layers to unleash the power of LLMs? How to bypass the hallucination effect of LLMs to meet stringent requirements for low-latency and ultra-reliability? Also, how to optimize wireless communications systems for efficient deployment and operation of LLMs in future networks?

### 09 Mar 2024 v4

@@ I Introduction
With the remarkable success of deep learning spanning from decision-making in AlphaGo to human-level interaction like ChatGPT, it is anticipated that artificial intelligence (AI) will be embodied in 6G networks. Along with the enhanced edge computing capabilities, AI could benefit the effective orchestration of network resources and improve the quality of service (QoS). Correspondingly, investigation on efficient AI-based service provisioning has attracted intense research interest. On the other hand, the application of one AI model is often limited to certain scenarios or tasks. In this context, large language models (LLMs) (e.g., generative pre-trained transformer, GPT) could perform well in various natural language processing (NLP) and computer vision tasks. These inspiring advancements shed light on revolutionizing cellular networks by LLMs. For instance, [1] harnesses collective intelligence for efficient network management, by delving into the deployment of on-device LLMs and proposing a multi-agent system architecture. Similarly, [2] challenges traditional network paradigms for LLM training, and proposes a novel, cost-effective architecture tailored to LLM-specific communication patterns, with a demonstrated $75\%$ network cost reduction without sacrificing the performance. These progresses [1, 2] underscore the importance of integrating LLMs with innovative network architectures, a key to unlock greater efficiency and performance in advanced network environments. Notably, towards provisioning personalized generative services (e.g., personalized assistance and recommendation systems), fine-tuning is still a prerequisite to align pre-trained LLMs to follow human intents [3] and yield personalized outputs. Nevertheless, it might be cost-ineffective to simply deploy multiple copies of bloated model parameters to support different purposes, and a feasible solution remains under-investigated.
@@ II Implementation Showcase of NetGPT
As illustrated in Fig. 2, we present a synergistic cloud-edge framework to accomplish personalized generative services, by leveraging distinctive pre-trained LLMs for cloud and edge (e.g., base stations [BSs]) deployment. In particular, limited by the availability of open-source LLMs, we select and deploy the LLaMA-7B model [8] and the GPT-2-base model, which consist of approximately $6.7$ and $0.1$ billion parameters, at the cloud and the edge, respectively. However, it should be noted that NetGPT allows the utilization of other LLMs as well. On this basis, we delve into implementation details of cloud-edge LLM synergy towards NetGPT in an incremental manner. In particular, we start with detailed DNN structures of two LLMs (i.e., LLaMA-7B model and GPT-2-base model). Then, we discuss the effective means to fine-tune these LLMs on computation-limited devices, and demonstrate the effectiveness of synergizing edge LLMs and cloud LLM for location-based personalized generative services. Notably, the “LLM synergy” framework significantly contrasts with split learning [5] and federated learning [3], which aims to train the divided segments of a DNN on different clients, or jointly learn from data distributed across multiple nodes in a data privacy-friendly manner. Orthogonal to split learning and federated learning, our framework focuses on effective cloud-edge collaboration with prompt enhancement & de-duplication at the edge and personalized responses at the cloud.
@@ II-B Low-Rank Adaptation and Cloud LLM Fine-Tuning
As LLaMA lacks the capability to generate responsive text [8], an extra fine-tuning is still required. However, a direct fine-tuning of LLMs such as a LLaMA still requires significant computational resources. For example, it demands $112$ GB video random access memory (VRAM) to fine-tune the LLaMA-7B model, far more than the capacity of NVIDIA A100 Tensor Core GPU. Therefore, we leverage a low-rank adaptation (LoRA) technique [11], a parameter efficient fine-tuning (PEFT) technique, to achieve parameter-efficient fine-tuning on a consumer-level hardware. Notably, for NetGPT, PEFT techniques are not indispensable components. Instead, LoRA just demonstrates the feasibility of fine-tuning on computation-limited network elements of NetGPT.
@@ II-C2 LLM-instructed data collection:
Different from the cloud LLM, which is fine-tuned with web-based conversational datasets [12], so to respond to “comprehensive prompts” in a manner that is informed, contextually rich, and aligns with the broader conversational context, the data collection for edge LLMs requires extra efforts. This differentiation in data sources underlines the collaborative efficacy of the synergy architecture, with each component playing a specialized role in the data processing workflow. In order to implement personalized edge LLM, it is crucial to grant the GPT-2-base model the capability to extend a “concise” prompt by appending location-based information. Basically, the positioning information can be conveniently obtained according to the locations of affiliated BSs stored in the 5G access and mobility management function (AMF). Meanwhile, in order to complement more comprehensive information, we take the self-instruct approach [13] and leverage OpenAI’s Text-Davinci-003 model to generate useful text samples. In particular, as for each location, we use a set of manually written location-related prompts to interact with the OpenAI’s Text-Davinci-003 model, and leverage the generative response texts as the “intended prompt”, which augments the most likely word following the “concise prompt” and describes the corresponding more comprehensive intent. For example, as illustrated on the top-left side of Fig. 3, from a perspective of real-life linguistic patterns, “libraries” frequently correlates with the contextual word “collections”. Besides, the top-right part of Fig. 3 demonstrates how post-fine-tuning edge LLM can generate an elongated and more comprehensive prompt in response to the “concise prompt”. On this basis, a series of mappings between the “concise prompt” and an “intended prompt” can be collected. Considering the size and task complexity of the edge LLM, we collect a dataset comprising approximately $4,000$ samples for directly fine-tuning the GPT-2-base model towards a prompt-completion model. Notably, for scenarios where stronger generality is required, edge LLMs can be enhanced with a larger-scale LLM, and fine-tuning techniques such as LoRA can be employed as well.
@@ II-E2 Comparison with cloud-only solution:
II-E3 Extension to large multi-modal models (LMMs)
The emergence of LMMs marks a significant leap beyond text-only tasks and NetGPT could naturally benefit from the integration with LMMs. For example, deploying smaller, versatile LMMs at the edge allows for the collection and refinement of local data, and generates prompts that are more comprehensively aligned with immediate surroundings. On the other hand, powerful LMMs located in the cloud, such as DALL·E and Sora, contribute to generating detailed multi-modal responses on top of enriched information from the edge. Notably, the effective integration with LMMs requires further attention due to multi-modal dataset scarcity, increased model complexity, and semantic gap across modalities.
@@ III AI-Native Network Architecture Towards NetGPT
We argue that NetGPT provides the opportunity to transform cellular networks into an AI-native networking architecture, which provisions personalized, networked and inclusive intelligence for end users and grants users more privilege to access generative services anytime and anywhere. Nevertheless, such a transformation does come at a cost. It requires substantial changes, far more than installing racks of servers at the edge location and local break-out of the traffic for edge processing. In particular, compared with conventional connectivity-oriented communications systems, wherein a typical service establishes connections between two specific terminals and the communication source and destination are clearly defined by end users, NetGPT requires to establish generative performance-driven connections more implicitly. Moreover, as NetGPT involves more frequent data collection and processing modules for training personalized LLM models, computing resources shall be consistently scheduled to accomplish NetGPT, and an independent computing plane (CmP), which coordinates computing resources and perform AI-related functionalities, becomes complementary to the control plane (CP) and user plane (UP). In other words, as shown in Fig. 4, NetGPT necessitates the design of deeply converged C&C in radio access networks (RANs). On top of these novel features, a logical AI workflow shall be devised to establish (beyond) generative service orchestration.
III-A Converged C&C Resource Management
As part of provisioned services in future cellular networks, the orchestration of resources for NetGPT shares some similarities with that for other network services, including seamless connectivity control in the CP and reliable information transmission in the UP. However, it also poses additional challenges. For example, for the advocated CmP in NetGPT, orchestrating heterogeneous computing resources effectively is of paramount importance, considering that the central processing units (CPUs) capably handle general computing tasks, while the graphics processing units (GPUs) specialize in parallel processing tasks such as neural network computations, and the video processing units (VPUs) excel in the real-time image and video processing. Therefore, it is crucial to accommodate tasks with the most appropriate resources to enhance system efficiency and performance, in an orchestrated manner with intelligent adaptation. In addition, since the scope of resources spans distributed nodes from the cloud to the terminal, novel protocol stack needs be carried on radio control signaling (e.g., RRC) or radio data protocol (e.g. PDCP or SDAP) to transmit AI-generative messages and implement model update & distribution. In that regard, introducing new RRC messages or flags that would prioritize AI data, could ensure rapid allocation of network resources for real-time AI workflows. Furthermore, as quantization techniques are widely used in developing light-weight AI models, refining PDCP header compression algorithms to better identify and compress AI data payloads promises improved data processing efficiency. Additionally, defining a dynamic configuration module/sublayer within the protocol stack, tailored to specific AI tasks, allows for the establishment of a closed-loop mechanism for real-time adjustments.
@@ III-C Personalized Profiling
III-D Logical AI Workflow
@@ V Conclusion & Future Work
  • How to evaluate the interpretability and reliability of the model’s outputs? How to bypass the hallucination effect of LLMs to meet stringent requirements for low-latency and ultra-reliability? More importantly, in addition to reinforcement learning from human feedback (RLHF), what techniques can be leveraged to satisfy users’ stringent requirements amidst significant model uncertainties, before practically deploying LLMs in scenarios where decision-making critically relies on AI insights?
  • How to incorporate the latest LMMs more effectively to provision personalized services beyond text? For different data modalities, how to effectively orchestrate heterogeneous computing resources to maximize operational efficiency?
  • Regarding the cloud-edge collaboration, how to accurately learn the variations of availability of computing resources and develop adaptive resource management strategies to dynamically respond to fluctuating computational demands? Besides personalized assistance and recommendation systems, other real-world applications (e.g., AI copilot, embodied AI agent, etc.) can be explored to fully unleash the capability of NetGPT. Correspondingly, how can NetGPT be evolved to meet the emerging demands?